At the beginning of the electronic tablet era, there was only one method to produce tablets. Summarizing inputs into labor L and capital K, the proportion K/L of capital to labor use in this method (also known as capital intensity) was fixed at 2 1. Derive the demand for labor L(w.r.q) as a function of the price of labor w, the price of capital r and of the output level q. [Warning: this function may be constant with respect to some variables.]
Answer #1

Production function where capital labor ratio is fixed at 2, is given by q = min(K, 2L). Now we see that K/L = 2 or K = 2L. Then we use q = min (2L, 2L). Hence the demand function for Labor is 2L = q or L = 0.5q. This is independent of input prices because labor and capital are used in fixed ratio.